OpenStudy (anonymous):

On a circle of radius 10, how long is an arc that subtends a central angle of 4pi/5 radians?

OpenStudy (unklerhaukus):

length of circumference is \(2 \pi r \) length of arc is \[\frac{2 \pi r}\theta \] where \(\theta\) is the angle the arc subtends